OpenStudy (anonymous):

how long is a guy wire reaching from the top of a 6ft pole, to a point on the ground, 4ft from the pole?

OpenStudy (anonymous):

Use Pythagorean Theorem: \[6^2 + 4^2 = x^2 \] \[36 + 16 = x^2\] \[52 = x^2 \] \[x \approx 7.211\] ft